### v3

# Not All Heads Matter: A Head-Level KV Cache Compression Method with Integrated Retrieval and Reasoning

Abstract
Key-Value (KV) caching is a common technique to enhance the computational efficiency of Large Language Models (LLMs), but its memory overhead grows rapidly with input length. Prior work has shown that not all tokens are equally important for text generation, proposing layer-level KV cache compression to selectively retain key information. Recognizing the distinct roles of attention heads in generation, we propose HeadKV, a head-level KV cache compression method, and HeadKV-R2, which leverages a novel contextual reasoning ability estimation for compression. Our approach operates at the level of individual heads, estimating their importance for contextual QA tasks that require both retrieval and reasoning capabilities. Extensive experiments across diverse benchmarks (LongBench, LooGLE), model architectures (e.g., Llama-3-8B-Instruct, Mistral-7B-Instruct), and long-context abilities tests demonstrate that our head-level KV cache compression significantly outperforms strong baselines, particularly in low-resource settings (KV size = 64 & 128). Notably, our method retains just 1.5% of the KV cache while achieving 97% of the performance of the full KV cache on the contextual question answering benchmark.  [^1] https://github.com/FYYFU/HeadKV/
1 Introduction
Modern Large Language Models (LLMs) increasingly support extremely long inputs: GPT-4 (Achiam et al., 2023), Llama-3 (Dubey et al., 2024), and Qwen-2 (Yang et al., 2024) handle up to 128K tokens, while Claude (Anthropic, 2024) supports up to 1 million tokens. These extended capacities improve performance on tasks like dialogue generation (Li et al., 2024a; Yi et al., 2024), question answering (Ho et al., 2020; Xu et al., 2023), and summarization (Xiao & Carenini, 2019; Koh et al., 2022). As input lengths increase, memory usage and latency grow significantly due to self-attention in transformers. To improve inference speed and efficiency, most LLM inference consists of two phases: prefilling for input processing and decoding for token generation, with key and value states from attention cached for reuse (KV cache). However, as input length increases, KV cache memory grows rapidly, posing significant challenges for storage and efficiency.
To address this, KV cache compression methods (Xiao et al., 2024; Li et al., 2024d; Cai et al., 2024; Feng et al., 2024) have been proposed, typically using token eviction to optimize retention per layer or head during prefilling, reducing memory without impacting performance. However, none have explored varying KV cache size across individual heads. Inspired by prior observations (Voita et al., 2019; Wu et al., 2024; Zheng et al., 2024) that attention heads vary in importance for generation, we propose HeadKV, a head-level KV cache compression method that allocates KV cache budgets based on head importance distribution using a novel retrieval and reasoning importance estimation. Specifically, heads deemed more important are allocated larger KV cache budgets, while less important ones receive smaller allocations, optimizing memory usage without sacrificing performance.
Figure 1: Our proposed head-level KV cache compression method consists of two steps: (1) Head-Level Importance Score Estimation (upper part): important heads that contribute to the contextual reasoning ability are identified using Needle-in-a-Haystack tests. (2) Head-Level KV Cache Allocation (lower part): KV cache budgets for each head during the prefilling phase are allocated based on the importance score distribution identified in the first step.
In addition to allocating KV cache across attention heads rather than layers, a key aspect of our approach is distributing cache budgets based on head importance measures. Prior work (Wu et al., 2024) proposed method to identify retrieval heads, using importance estimation to assess each head’s role in retrieving relevant context. We integrate their measure with our head-level KV cache compression as a baseline, observing improved performance over layer-level cache compression.
However, we argue that allocating larger KV cache budgets solely to retrieval heads is insufficient for tasks like contextual Question Answering (QA), which requires both retrieval and reasoning to handle long input contexts effectively. To address this, we propose an importance-score estimation method that jointly evaluates each head’s retrieval and reasoning capabilities for KV cache allocation. Specifically, as illustrated in the Importance Score Estimation section of Figure 1, we construct questions that require both the retrieval and reasoning abilities of the LLM. For instance, in the provided example, the model must first identify “who is younger between John and Mary” (John) by referencing the context of their ages, and then retrieve John’s favorite thing. We then estimate the importance score of each head based on the attention scores generated by the model while answering the question. Using the estimated importance scores, we allocate the KV cache budget for each individual head, meaning that heads demonstrating greater importance in retrieval and reasoning retain a larger portion of the KV cache, as shown in the Head-Level Allocation section of Figure 1. Within each head, we then retain only the most relevant KV cache entries, following the strategy proposed in Li et al. (2024d).
We conduct experiments on various benchmarks requiring both retrieval and reasoning abilities, including QA tasks from LongBench (Bai et al., 2024) and LooGLE (Li et al., 2024b), using backbone models such as Llama-3-8B-Instruct (Dubey et al., 2024) and Mistral-7B-Instruct (Jiang et al., 2023). Our results demonstrate that the head-level KV cache compression method outperforms previous approaches on nearly all tasks. Furthermore, the allocation strategy based on our estimated importance scores—reflecting both retrieval and reasoning abilities—outperforms allocation strategies based on retrieval ability alone. In challenging scenarios like needle-in-a-Haystack and reasoning-in-a-Haystack tests, our methods effectively preserve the model’s retrieval and reasoning capabilities. Finally, experiments on memory and latency reveal that our approach significantly reduces memory usage and decoding latency, comparing with the original full KV.
2 Related Work
2.1 Attention Heads
Multi-Head Attention, a fundamental component of Transformer architectures (Vaswani et al., 2017), has been extensively analyzed to understand the roles of individual attention heads (Voita et al., 2019; Olsson et al., 2022; Jin et al., 2024; Wu et al., 2024; Zheng et al., 2024), often with goals such as pruning redundant heads(Shim et al., 2021; Kwon et al., 2022) or enhancing interpretability(Olsson et al., 2022; Jin et al., 2024; Zheng et al., 2024). For example, Voita et al. (2019) observed that only a small subset of heads plays a key role in machine translation, typically managing positional information, syntactic parsing, or focusing on rare words. Similarly, Olsson et al. (2022) identified ’induction heads’ that implement a copy-and-paste mechanism, enabling models to replicate previously encountered sequences. Additionally, Zheng et al. (2024) provided a thorough overview of recent efforts to characterize the diverse functions of attention heads, categorizing them into four types: Knowledge Recalling, In-Context Identification, Latent Reasoning, and Expression Preparation. A closely related study by Wu et al. (2024) discovered “retrieval heads” that play a crucial role in knowledge acquisition, using Needle-in-a-Haystack tests. These insights into head-level functionality serve as the foundation for our head-level KV cache compression methods, designed to jointly preserve both retrieval and reasoning capabilities during the compression process.
2.2 KV cache compression
Improving computational efficiency for LLMs, particularly for extremely long inputs, has attracted considerable research interest (Shazeer, 2019; Chen et al., 2023; Dao, 2024; Zhou et al., 2024; Ye et al., 2024), including caching key and value vectors in Multi-Head Attention to improve generation efficiency (Xiao et al., 2024; Zhang et al., 2023; Li et al., 2024d; Cai et al., 2024). One challenge with full KV cache approaches is that the memory usage increases linearly with input length, leading to significant memory management challenges. Various KV cache compression techniques have been proposed to reduce memory usage and improve inference efficiency. For example, Xiao et al. (2024) addressed the ‘attention sink’ issue with StreamingLLM, retaining only the first $k$ tokens’ KV cache. Zhang et al. (2023) applied the Heavy Hitter Oracle strategy to select key cache entries, while Li et al. (2024d) used the attention scores of the last $\alpha$ tokens to identify relevant entries. Cai et al. (2024) introduced PyramidKV, assigning smaller cache budgets to higher layers based on attention matrix patterns.
While these methods have improved performance and efficiency, they largely depend on rigid layer-level allocation strategies, which may not fully optimize KV cache allocation for downstream tasks. Feng et al. (2024) proposed dynamic head-level allocation using attention scores but still relied on layer-level budgeting. Similarly, Tang et al. (2024) employed Retrieval Heads distribution for head-level allocation but retained a Full KV cache for key heads, limiting true head-level compression. In contrast, our approach allocates KV cache solely based on head-level importance scores, independent of layer constraints, resulting in more effective compression and outperforming baselines without increasing decoding latency or memory usage.
3 Method
This section presents our proposed head-level KV cache compression method, which consists of three key components: (1) identifying important heads and calculating head-level importance score distributions (3.1), (2) using these distributions to efficiently allocate KV cache budgets across heads (3.2), and (3) determining which Key and Value vectors to retain within each head (3.3).
3.1 Head-level Importance Score Estimation
Accurate budget allocation at the head level requires identifying which heads are most and least important for the given task. By leveraging this importance distribution, we can assign larger KV cache budgets to more critical heads and smaller budgets to less significant ones. To achieve this, we propose a novel importance-score estimation method, inspired by Wu et al. (2024), which allows us to effectively estimate the importance of each attention head for optimal KV cache allocation.
Figure 2: Comparison of examples for head identification: Needle-in-a-Haystack test example from Wu et al. (2024) for identifying Retrieval Heads distribution (left), and our proposed Needle-in-a-Haystack test example for identifying Retrieval-Reasoning Heads distribution (right).
Retrieval Heads
Wu et al. (2024) use the Needle-in-a-Haystack test  [^2] https://github.com/gkamradt/LLMTest_NeedleInAHaystack and custom retrieval examples, as shown in Figure 2, to estimate the importance score for each head. In these examples, a question ${\bm{q}}$ that cannot be answered using the model’s parametric knowledge is paired with an answer ${\bm{k}}$ (the “Needle”) inserted into a haystack ${\bm{c}}$ at different positions $p_{i}$. The model is required to retrieve the exact answer ${\bm{k}}$ from the combined input $({\bm{k}},{\bm{c}})$.
During each decoding step $t$, a head $h$ earns a fraction of the importance score if (1) the token with the highest attention score $\arg\max({\bm{a}}_{h})$ matches the generated token, and (2) the token is part of the inserted answer ${\bm{k}}$. The final importance score for each head $h$ is calculated accordingly:
$S_{h}=\sum_{t=1}^{N}\mathcal{N}^{t},\quad\text{where}\ \mathcal{N}^{t}=\begin{cases}\frac{1}{N}&\text{if }\arg max({\bm{a}}_{h})\in{\bm{k}}\\ 0&\text{otherwise}\\ \end{cases}$ (1)
where $N$ is the length of the inserted answer ${\bm{k}}$, and ${\bm{a}}_{h}$ is the attention score on the combined input from head $h$. By using various settings of the inserted position $p_{i}$, they obtain the head-level Retrieval Heads distribution.
Directly using this distribution poses two issues: (1) it focuses on the retrieval-and-paste mechanism, lacking consideration for the contextual and reasoning skills needed for complex questions; and (2) the distribution is too sparse for effective head-level KV cache allocation, with nearly 70% of heads receiving an importance score of zero due to the strict exact-match requirement (Wu et al., 2024).
Retrieval-Reasoning (R2) Heads
To address these issues, we propose a new importance score estimation method that accounts for both the retrieval and reasoning abilities of the heads, enabling a more accurate assessment of their significance.
First, we construct retrieval-reasoning examples by adding explicit contextual reasoning steps to the retrieval examples from Wu et al. (2024), as shown in the Retrieval-Reasoning Example part of Figure 2. We further modify the inserted needle into three parts: ${\bm{k}}=({\bm{r}},{\bm{c}}^{1},{\bm{c}}^{2})$, where ${\bm{r}}$ is the reasoning step, and ${\bm{c}}^{1}$ and ${\bm{c}}^{2}$ are different answers to the refined question ${\bm{q}}$. The model must reason with ${\bm{r}}$ to retrieve and generate the correct answer ${\bm{c}}^{2}$, avoiding the wrong answer ${\bm{c}}^{1}$.
Secondly, we refine the estimation method by focusing on the entire correct answer ${\bm{c}}^{2}$ (Correct Copy in Figure 1), since all tokens are relevant to the question ${\bm{q}}$. This approach aligns with Contextual QA, which requires both retrieval and reasoning abilities. By considering the full correct answer, the importance score for each head $h$ no longer depends solely on the token with the highest attention score. The importance score for head $h$ is calculated as follows:
$S_{h}=\sum_{t=1}^{N}\sum_{i=1}^{N}\mathcal{M}^{t}_{i},\quad\text{where}\ \mathcal{M}^{t}_{i}=\begin{cases}\frac{{\bm{a}}_{i}}{N}&\text{if }\text{top-}i({\bm{a}}_{h})\in{\bm{c}}^{2}\\ 0&\text{otherwise}\\ \end{cases}$ (2)
[…]
3.2 Head-Level KV Cache Allocation
With the importance scores estimated for each head, we can identify the key heads and allocate the KV cache budget accordingly. In this section, we explain how to incorporate these distributions into head-level KV cache allocation.
Preliminary
In Multi-Head Attention, for each head $h$ in layer $l$, the embedded input ${\bm{X}}=\{x_{1},x_{2},\dots,x_{n}\}\in\mathbb{R}^{n\times d}$ is mapped into different subspaces using the query ${\bm{W}}^{l}_{Q}$, key ${\bm{W}}^{l}_{K}$, and value ${\bm{W}}^{l}_{V}\in\mathbb{R}^{d\times d_{h}}$ matrices:
${\bm{Q}}^{l}_{h}={\bm{X}}{\bm{W}}^{l}_{Q}\in\mathbb{R}^{n\times d_{h}};\quad{\bm{K}}^{l}_{h}={\bm{X}}{\bm{W}}^{l}_{K}\in\mathbb{R}^{n\times d_{h}};\quad{\bm{V}}^{l}_{h}={\bm{X}}{\bm{W}}^{l}_{V}\in\mathbb{R}^{n\times d_{h}}.$ (3)
To optimize memory and enhance efficiency, KV cache compression methods (Xiao et al., 2024; Li et al., 2024d; Cai et al., 2024) are employed to discard unimportant KV cache entries while preserving performance. For each head $h$, the compressed KV cache is reduced to ${\bm{K}}^{l}_{h}\in\mathbb{R}^{s\times d_{h}}$ and ${\bm{V}}^{l}_{h}\in\mathbb{R}^{s\times d_{h}}$, where $s\ll n$, resulting in a significant improvement to computational efficiency.
Head-level Allocation
Previous works on KV cache compression during the prefill phase (Xiao et al., 2024; Li et al., 2024d; Cai et al., 2024; Feng et al., 2024) are limited to layer-level allocation, using either uniform or dynamic budgets per layer, but treating all heads within a layer equally. While Feng et al. (2024) incorporate head-level information, their approach still depends on layer-level allocation as a prerequisite.
Building on the head-level importance distributions, we propose a comprehensive KV cache allocation strategy. Each head $h$ is initially assigned a fixed KV cache size $b$ with an associated importance score $S_{h}$. To allow dynamic allocation, we create a shared budget pool $B$ by extracting a portion of the budget from each head, leaving the remainder as the basic budget. This process is illustrated in the Head-Level Allocation section of Figure 1. The budget pool $B$ is then distributed among the heads in proportion to their importance scores $S_{h}$. The final head-level KV cache allocation is as follows:
$B=\frac{b}{\beta}\times L\times H;\quad{\bm{b}}_{h}=(b-\frac{b}{\beta})+S_{h}\times B$ (4)
where $b$ is the initial fix budget for each head, $\beta$ is a hyper-parameter to control the size of the dynamic budget pool, $L$ and $H$ is the numbers of layers and heads of current LLM respectively.
The last $\alpha$ instruct tokens are preserved before forming the dynamic budget pool $B$ to guide the selection process, as detailed in Section 3.3. The retained KV cache for each head includes: (1) the basic budget ($b-\frac{b}{\beta}$), (2) the dynamic budget $S_{h}\times B$, proportional to its importance score, and (3) the last $\alpha$ instruct tokens.
3.3 KV Cache Selection
After determining the number of KV cache entries to retain using the above algorithm, we apply an attention-based selection strategy from prior works (Li et al., 2024d; Cai et al., 2024; Feng et al., 2024) to keep the most relevant entries. Specifically, the last $\alpha$ instruction tokens (local windows) guide KV cache selection for each head. Attention scores from these local windows to the remaining tokens are aggregated through a pooling layer, with higher-scoring tokens considered more important and retained in the cache.
4 Experiments And Analysis
This section outlines the experimental setup, including KV cache baselines and implementation details. We also conduct additional experiments that: (1) emphasize the importance of enhancing contextual reasoning in importance score estimation (4.3); (2) use the Needle-in-a-Haystack and Reasoning-in-a-Haystack tests to demonstrate how our head-level KV cache compression improves long-context retrieval and reasoning (4.4); and (3) provide a comprehensive comparison with previous methods, showing our approach delivers superior performance while maintaining computational efficiency (4.5).
4.1 Experiment settings
Models and Datasets
We compare our head-level KV cache compression method against strong baselines using two open-source LLMs: Llama-3-8B-Instruct (Dubey et al., 2024) and Mistral-7B-Instruct (Jiang et al., 2023). The evaluation is based on two benchmarks for long-context understanding: LongBench (Bai et al., 2024) and LooGLE (Li et al., 2024b). For LongBench, we use datasets from the Single-Doc QA and Multi-Doc QA categories to assess contextual reasoning. For LooGLE, we focus on the Long Dependency QA task, which includes four QA-related tasks. Dataset details are in Appendix Table 5.
Baselines and Settings
We evaluate three strong KV cache compression methods as baselines, ensuring all retain the same number of KV cache entries for fair comparison:
  • 1)
    SnapKV (Li et al., 2024d) uses the last $\alpha$ tokens as local windows to guide KV cache selection. Attention scores from these windows to the remaining tokens are pooled to cluster information and guide the selection process.
  • 2)
    PyramidKV (Cai et al., 2024) follows a pyramid attention pattern, allocating more KV cache to lower layers to retain key information, while reducing the budget for higher layers where information is already aggregated.
  • 3)
    Ada-KV (Feng et al., 2024) dynamically allocates budgets to heads within each layer based on their concentration degrees, and can be combined with SnapKV or PyramidKV. Ada-SnapKV is used as the baseline due to its superior performance over Ada-PyramidKV.
Our proposed head-level KV cache compression method also requires a strategy to guide KV cache selection after allocating the budget. Therefore, we use the SnapKV method to guide the selection process for each head. We set the size of the local windows $\alpha=8$ for both the baselines and our method. The hyper-parameter $\beta$, which controls the size of the shared budget pool, was chosen from {1.005, 1.01, 1.1, 1.2, 1.5, 2, 5, 10}, and we report the best performance.
4.2 Main Results
Method Single-Doc QA Multi-Doc QA Avg. Long dependency QA Avg.
NartvQA Qasper MF-en HotpotQA 2WikiMQA Musique Doc.QA Info. Retrieval Timeline Computation
Llama-3-8B-Instruct, KV Size = Full
FullKV 25.56 32.07 39.71 43.57 35.28 21.18 32.90 8.73 11.21 0.67 7.43 7.01
Llama-3-8B-Instruct, KV Size = 128
SnapKV 22.11 15.79 31.01 41.12 29.20 19.35 26.43 8.36 9.46 0.79 6.56 6.29
PyramidKV 22.01 17.05 31.52 39.27 28.99 18.34 26.20 8.89 9.63 0.61 6.72 6.46
Ada-SnapKV 22.99 19.95 34.22 42.97 30.82 20.15 28.52 9.07 10.3 0.54 6.59 6.63
HeadKV-R 23.49 25.39 38.15 42.45 32.84 19.95 30.38 8.87 10.35 0.78 7.52 6.88
HeadKV-R2 21.80 29.19 41.89 43.73 35.01 20.40 32.00 9.60 11.13 0.67 7.22 7.16
Llama-3-8B-Instruct, KV Size = 1024
SnapKV 25.76 27.50 38.38 43.40 34.81 20.07 31.65 9.61 11.34 0.53 7.22 7.18
PyramidKV 25.38 26.83 36.90 44.09 34.24 21.49 31.49 8.98 11.41 0.53 6.96 6.97
Ada-SnapKV 25.79 29.24 38.74 43.93 36.34 19.79 32.31 8.65 11.41 0.53 7.71 7.08
HeadKV-R 24.85 30.94 39.82 43.52 36.58 20.37 32.68 9.20 11.67 0.55 7.71 7.28
HeadKV-R2 24.66 30.82 39.56 43.97 36.47 22.24 32.95 9.02 11.51 0.47 7.85 7.21
Mistral-7B-Instruct, KV Size = Full
FullKV 26.63 32.99 49.34 42.77 27.35 18.78 32.98 12.17 15.52 0.49 10.03 9.55
Mistral-7B-Instruct, KV Size = 128
SnapKV 21.47 21.95 45.24 33.88 21.83 15.53 26.65 10.86 12.24 0.57 8.81 8.12
PyramidKV 21.76 21.98 43.72 32.76 22.73 15.59 26.42 10.64 11.9 0.47 8.69 7.93
Ada-SnapKV 21.57 24.59 46.70 35.74 25.57 14.37 28.09 11.14 12.37 0.45 9.57 8.38
HeadKV-R 23.97 29.60 48.40 39.66 26.31 18.13 31.01 11.43 13.04 0.53 10.26 8.82
HeadKV-R2 25.04 27.95 48.48 41.28 27.65 18.05 31.41 11.44 13.08 0.63 10.20 8.84
Mistral-7B-Instruct, KV Size = 1024
SnapKV 25.38 30.22 49.29 41.84 26.60 18.08 31.90 11.69 13.89 0.52 10.54 9.16
PyramidKV 24.28 30.05 49.17 40.49 26.43 18.80 31.54 11.77 14.51 0.51 10.19 9.25
Ada-SnapKV 24.82 31.49 48.80 41.18 27.38 18.22 31.98 11.96 13.82 0.53 9.92 9.06
HeadKV-R 25.87 31.44 49.55 41.95 27.09 19.88 32.63 12.21 14.17 0.50 10.58 9.37
HeadKV-R2 25.64 32.54 50.49 41.80 27.88 18.89 32.87 11.94 14.93 0.50 10.49 9.47
Table 1: Performance comparison on the LongBench and LooGLE benchmarks for Llama-3-8B-Instruct and Mistral-7B-Instruct. Our head-level KV cache compression method outperforms all baselines, especially in low-resource settings (KV size = 128). It even exceeds the FullKV result (32.90) on Llama-3-8B-Instruct (KV size = 1024, 32.95), highlighting the benefits of incorporating contextual reasoning for head selection.
Table 1 lists the evaluation results for contextual tasks from the LongBench and LooGLE benchmarks. Our head-level KV cache compression method consistently outperforms strong baselines, especially with 64 and 128 KV cache configurations. In resource-constrained settings, precise KV cache allocation is crucial. Layer-level methods allocate a fixed cache size to all heads within a layer, making it difficult to retain essential information. Ada-SnapKV improves this by allowing dynamic allocation within layers, but still relies on fixed layer-level budgets. In contrast, our head-level strategy allocates dynamically across individual heads, retaining critical information by adjusting the budget based on each head’s importance.
We perform head-level allocation using both the standard Retrieval Heads distribution (HeadKV-R) and our Retrieval-Reasoning Heads distribution (HeadKV-R2) for global KV cache allocation. This combination leads to superior performance across benchmarks. Notably, integrating the Retrieval-Reasoning Heads distribution significantly improves results over the standard Retrieval Heads distribution, highlighting the impact of our approach. Our Retrieval-Reasoning distribution even surpasses the Full-KV cache average (32.90), achieving 32.95 with a 1024 KV cache. Overall, both the head-level allocation strategy and Retrieval-Reasoning distribution are key to these performance gains.
In addition, we present the results for various retained KV cache sizes (64, 128, 256, 512, 1024) in Figure 3, , with detailed results available in Appendix Table 4.
Figure 3: Results for different KV cache sizes (64, 128, 256, 512, 1024), showing average accuracy across six datasets from the LongBench benchmark with an average input length of 8,683 tokens. Notably, a KV cache size of 64 retains just 0.7% of the total tokens.
4.3 Retrieval-Reasoning Heads
As detailed in Section 3.1, we propose to improve the standard Retrieval Heads distribution by incorporating retrieval-reasoning examples and refining importance score estimation to better capture contextual reasoning and identify relevant heads. We also conduct an ablation study to evaluate the impact of these modifications.
Figure 4: Head visualization for Llama-3-8B-Instruct results. The Retrieval Heads distribution is sparse to effectively differentiate between heads, while our Retrieval-Reasoning Heads has denser distribution for such differentiation. See Appendix Figure7 for Mistral-7B-Instruct results.
Method Single-Doc QA Multi-Doc QA Avg.
NartvQA Qasper MF-en HotpotQA 2WikiMQA Musique
Llama-3-8B-Instruct, KV Size=128
HeadKV-R 23.49 25.39 38.15 42.45 32.84 19.95 30.38
HeadKV-ER 23.33 25.86 40.28 43.25 33.23 20.28 31.04
HeadKV-R2 21.80 29.19 41.89 43.73 35.01 20.40 32.00
Mistral-7B-Instruct, KV Size=128
HeadKV-R 23.97 29.60 48.40 39.66 26.31 18.13 31.01
HeadKV-ER 23.23 28.70 48.10 41.39 27.31 17.39 31.02
HeadKV-R2 25.04 27.95 48.48 41.28 27.65 18.05 31.41
Table 2: Ablation study results on the LongBench benchmarks. HeadKV-R leverages the standard Retrieval Heads distribution. HeadKV-ER uses the retrieval examples from Wu et al. (2024) but with our proposed importance score estimation method. HeadKV-R2 leverages both our proposed importance score estimation method and the retrieval-reasoning examples.
Table 2 presents the ablation study results, while Figure 4 provides visualizations for each distribution. Alongside the standard Retrieval Heads and Retrieval-Reasoning Heads distributions, we introduce the Enhanced-Retrieval Heads distribution, using retrieval examples with our modified importance score estimation method. Comparing Retrieval Heads (HeadKV-R) and Enhanced-Retrieval Heads (HeadKV-ER) reveals that focusing on the entire needle, rather than specific tokens, improves performance. Figure 4 shows that the Retrieval Heads distribution is sparse, while the Enhanced-Retrieval and Retrieval-Reasoning distributions are much denser. The strict constraints on the Retrieval Heads distribution result in most heads receiving a score of zero, leading to a worse results when incorporating Retrieval Heads distributions.
While the Enhanced-Retrieval Heads distribution improves performance slightly, it remains rely on retrieval examples and lacks full consideration of contextual reasoning. In contrast, the Retrieval-Reasoning Heads distribution, reflecting both retrieval and reasoning abilities, consistently outperforms other methods, underscoring the value of incorporating retrieval-reasoning examples.
4.4 Long-Context Retrieval and Reasoning
Figure 5: Neele-in-a-Haystack test results on Llama-3-8B-Instruct with KV cache = 128. We build our head-level KV cache method based on SnapKV and our proposed method significantly outperform all strong baselines. Moreover, our Retrieval-Reasoning Heads distribution maintains and improves long context retrieval ability. Results on Mistral-7B-Instruct can be found in Appendix Figure 8, which are consistent with results on Llama-3-8B-Instruct.
We conduct the Needle-in-a-Haystack test to assess the long-context retrieval ability of different KV cache compression methods. As illustrated in Figure 5, we set the KV size to 128 for all methods and keep the other hyperparameters consistent with previous experiments. Results from the Llama-3-8B-Instruct demonstrate that our head-level KV cache compression method effectively retain important information compared with other strong baselines, verifying the superior long-context retrieval ability of our proposed method. However, these tests only retrieve the inserted needle from the haystack and paste it into the generation, lacking an assessment of contextual reasoning ability in long-context scenarios. This long-context contextual reasoning ability is crucial for many tasks, such as question answering (QA), summarization (Kuratov et al., 2024; Li et al., 2024c). Therefore, we conduct the Reasoning-in-a-Haystack test to evaluate the long-context reasoning ability of each KV cache compression method across different scenarios.
We follow the setup from Kuratov et al. (2024) to conduct the Reasoning-in-a-Haystack test. This test enhances the bAbI benchmark (Weston et al., 2015), designed for reasoning evaluation, by using text from the PG19 dataset (Rae et al., 2020) as the haystack. Reasoning needles from bAbI are inserted into the haystack, and the model must retrieve and reason through them to generate the correct answer. We use the dataset from Kuratov et al. (2024), averaging results across QA1-QA5 tasks for evaluation. Examples of tasks are shown in Figure 9.
As shown in Table 3, our head-level KV cache compression method significantly outperforms strong baselines, demonstrating its superior long-context reasoning. By incorporating retrieval-reasoning examples, our method achieves even better accuracy, particularly with the Llama-3-8B-Instruct model. Notably, combining head-level KV cache allocation with the standard Retrieval Heads distribution also yields improved results over other baselines. This is due to two factors: first, as shown in Figure 4, there is overlap between Retrieval and Retrieval-Reasoning Heads, indicating heads may serve multiple roles. Second, since the bAbI benchmark contains the answer within the inserted needle (see Figure 9), emphasizing retrieval alone helps our method locate the needle.
Method 0k 1k 2k 4k 8k Avg. 0k 1k 2k 4k 8k 16k 32k Avg.
FullKV 63.00 59.40 55.80 56.60 50.40 57.04 61.40 55.20 52.40 40.80 40.40 35.00 31.80 45.29
Llama-3-8B-Instruct, KV Size=128 Mistral-7B-Instruct, KV Size=128
SnapKV 60.80 57.00 54.80 52.60 45.60 54.16 56.60 50.60 46.60 34.40 35.40 33.20 29.00 40.83
PyramidKV 61.20 58.00 52.80 52.60 47.60 54.44 58.40 50.60 46.80 36.00 36.20 31.20 28.40 41.09
Ada-SnapKV 61.80 59.20 53.80 53.60 46.60 55.00 58.00 51.20 46.20 35.40 36.40 31.60 28.80 41.09
HeadKV-R 61.60 57.00 52.60 55.00 49.60 55.16 58.40 54.00 50.80 38.00 37.40 31.80 30.20 42.94
HeadKV-R2 62.60 60.40 55.00 55.80 50.40 56.84 60.20 54.20 51.20 37.20 37.40 32.60 28.80 43.09
Table 3: Reasoning-in-a-Haystack test results with KV cache = 128. The final results are averaged across QA1-QA5 tasks for each length. Unlike the Needle-in-a-Haystack test, this test inserts multiple needles into the haystack, requiring the model to reason through them to extract the correct answer.
4.5 Memory & Latency
Figure 6: The Decoding Latency and Peak Memory Usage results. Our proposed method maintain the comparable computational efficiency with other KV cache compression baselines.
In this section, we evaluate the computational efficiency of our head-level KV cache compression method using the Mistral-7B-Instruct model with a maximum sequence length of 32K, and FlashAttention (Dao, 2024) as the default setting.
To assess the decoding latency of each method, we use 32K-length data from the Reasoning-in-a-Haystack test as input and set various generation lengths (1, 512, 1024, 2048, 4096) for comparison. As shown in the Decoding Latency of Figure 6, our proposed method achieves the same decoding latency as other KV cache compression methods while maintaining performance closest to the Full KV cache. Notably, the decoding time includes both the pre-filling time and the decoding time. Therefore, we can conclude that the pre-filling time for our method and other baselines (such as PyramidKV, Ada-SnapKV) is almost negligible, as shown at the starting point (generation length = 1) of Decoding Latency in Figure 6.
In addition to decoding latency, KV cache compression methods also aim to reduce memory usage during the decoding phase. Therefore, we provide the Peak Memory Usage results, as shown in the Peak Memory Usage of Figure 6. All results for Peak Memory Usage are averaged over three trials. Our proposed methods achieve performance comparable to other KV cache compression baselines, significantly reducing memory usage compared to the Full KV cache.
5 Conclusion and Future Work
In this paper, we introduce HeadKV-R2, a head-level KV cache compression method with two key components. We propose a strategy for allocating KV cache budgets across attention heads based on their importance scores and develop a novel approach to estimate these scores by constructing examples that account for both retrieval and reasoning abilities. By allocating larger KV cache budgets to more important heads and smaller budgets to less important ones, our approach efficiently retain important KV cache than layer-level KV cache compression methods. We thoroughly evaluate HeadKV across multiple benchmarks, models, and long-context ability tests, with comprehensive results demonstrating that our method achieves superior performance while maintaining computational efficiency.
For future work, further exploration of various types of attention heads could offer valuable insights, such as those involved in in-context learning (Olsson et al., 2022) and truthfulness (Li et al., 2023). For example, placing greater emphasis on truthfulness heads could help mitigate issues like hallucination, thereby improving the factual accuracy of model outputs. Additionally, it would be worthwhile to investigate the development of a general task-specific score estimation algorithm. One potential direction is to leverage gradients from specific tasks to allocate KV cache budgets more effectively, enabling tailored compression to enhance model performance across diverse applications.
Appendix A Detaile Results
In Table 4, we show the detailed results of Figure 3 in the main paper.
Method Single-Doc QA Multi-Doc QA Avg. Long dependency QA Avg.
NartvQA Qasper MF-en HotpotQA 2WikiMQA Musique Doc.QA Info. Retrieval Timeline Computation
Llama-3-8B-Instruct, KV Size=Full
FKV 25.56 32.07 39.71 43.57 35.28 21.18 32.90 8.73 11.21 0.67 7.43 7.01
Llama-3-8B-Instruct, KV Size=64
SKV 20.51 12.80 31.69 37.02 25.91 17.02 24.16 8.84 9.43 0.66 6.18 6.28
PyramidKV 21.17 13.66 29.34 34.86 23.46 15.88 23.06 8.27 9.31 0.63 6.86 6.27
Ada-SKV 22.26 17.30 33.37 39.82 27.86 17.85 26.41 9.08 9.86 0.55 6.82 6.58
Ours. Copy 22.67 23.54 37.51 37.45 29.76 19.01 28.32 8.80 10.51 0.58 6.68 6.64
Ours. Mix 23.21 25.33 38.71 40.64 31.33 19.35 29.76 9.46 10.66 0.61 6.92 6.91
Llama-3-8B-Instruct, KV Size=128
SKV 22.11 15.79 31.01 41.12 29.20 19.35 26.43 8.36 9.46 0.79 6.56 6.29
PyramidKV 22.01 17.05 31.52 39.27 28.99 18.34 26.20 8.89 9.63 0.61 6.72 6.46
Ada-SKV 22.99 19.95 34.22 42.97 30.82 20.15 28.52 9.07 10.3 0.54 6.59 6.63
Ours. Copy 23.49 25.39 38.15 42.45 32.84 19.95 30.38 8.87 10.35 0.78 7.52 6.88
Ours. Mix 21.80 29.19 41.89 43.73 35.01 20.40 32.00 9.60 11.13 0.67 7.22 7.16
Llama-3-8B-Instruct, KV Size=256
SKV 23.38 20.18 37.65 42.80 33.23 20.01 29.54 9.04 10.59 0.53 7.53 6.92
PyramidKV 23.94 20.27 36.27 42.51 31.44 19.99 29.07 8.66 10.61 0.53 6.98 6.70
Ada-SKV 24.20 24.63 37.95 43.64 33.27 20.03 30.62 9.29 11.23 0.62 7.10 7.06
Ours. Copy 23.83 29.04 39.90 42.36 33.58 20.57 31.54 9.05 11.15 0.52 7.22 6.99
Ours. Mix 24.68 30.49 38.59 44.32 36.41 20.54 32.51 9.47 11.56 0.54 7.65 7.31
Llama-3-8B-Instruct, KV Size=512
SKV 25.47 23.75 38.64 43.66 33.98 19.83 30.89 9.00 11.07 0.63 7.34 7.01
PyramidKV 24.69 23.65 35.10 43.25 31.16 20.06 29.65 8.90 10.62 0.74 7.57 6.96
Ada-SKV 25.73 25.44 37.84 43.78 33.95 19.83 31.09 9.22 11.18 0.53 7.42 7.09
Ours. Copy 23.84 29.21 39.79 44.41 36.09 20.59 32.32 9.13 11.61 0.56 7.12 7.11
Ours. Mix 24.75 29.75 38.03 44.43 36.45 21.67 32.51 9.34 11.26 0.56 7.54 7.18
Llama-3-8B-Instruct, KV Size=1024
SKV 25.76 27.50 38.38 43.40 34.81 20.07 31.65 9.61 11.34 0.53 7.22 7.18
PyramidKV 25.38 26.83 36.90 44.09 34.24 21.49 31.49 8.98 11.41 0.53 6.96 6.97
Ada-SKV 25.79 29.24 38.74 43.93 36.34 19.79 32.31 8.65 11.41 0.53 7.71 7.08
Ours. Copy 24.85 30.94 39.82 43.52 36.58 20.37 32.68 9.20 11.67 0.55 7.71 7.28
Ours. Mix 24.66 30.82 39.56 43.97 36.47 22.24 32.95 9.02 11.51 0.47 7.85 7.21
Mistral-7B-Instruct, KV Size=Full
FKV 26.63 32.99 49.34 42.77 27.35 18.78 32.98 12.17 15.52 0.49 10.03 9.55
Mistral-7B-Instruct, KV Size=64
SKV 19.95 18.63 38.16 31.24 21.39 13.81 23.86 10.41 11.49 0.46 9.38 7.94
PyramidKV 20.91 19.61 38.05 32.18 22.87 15.26 24.81 10.64 11.69 0.56 9.06 7.99
Ada-SKV 22.70 21.28 42.39 34.35 22.40 14.85 26.33 10.56 11.50 0.45 8.72 7.81
Ours. Copy 24.23 25.22 46.02 38.82 26.05 17.41 29.63 10.94 13.14 0.63 9.11 8.46
Ours. Mix 21.77 26.57 48.39 40.12 26.76 16.21 29.97 11.19 13.94 0.48 9.87 8.87
Mistral-7B-Instruct, KV Size=128
SKV 21.47 21.95 45.24 33.88 21.83 15.53 26.65 10.86 12.24 0.57 8.81 8.12
PyramidKV 21.76 21.98 43.72 32.76 22.73 15.59 26.42 10.64 11.9 0.47 8.69 7.93
Ada-SKV 21.57 24.59 46.70 35.74 25.57 14.37 28.09 11.14 12.37 0.45 9.57 8.38
Ours. Copy 23.97 29.60 48.40 39.66 26.31 18.13 31.01 11.43 13.04 0.53 10.26 8.82
Ours. Mix 25.04 27.95 48.48 41.28 27.65 18.05 31.41 11.44 13.08 0.63 10.20 8.84
Mistral-7B-Instruct, KV Size=256
SKV 22.26 24.94 48.30 36.76 25.16 14.93 28.72 11.07 12.39 0.53 9.13 8.28
PyramidKV 21.42 25.36 47.94 38.75 25.82 15.30 29.10 11.57 12.35 0.56 9.51 8.50
Ada-SKV 23.81 27.04 48.56 38.32 25.34 16.42 29.91 11.67 13.57 0.52 10.53 9.07
Ours. Copy 24.98 29.31 49.01 41.36 27.16 17.34 31.53 11.94 13.30 0.63 10.95 9.21
Ours. Mix 24.94 31.02 50.76 42.11 26.14 18.47 32.24 12.37 13.88 0.48 9.86 9.15
Mistral-7B-Instruct, KV Size=512
SKV 24.18 28.87 48.74 38.84 25.48 15.04 30.19 11.96 13.47 0.52 10.50 9.11
PyramidKV 23.07 28.97 48.37 39.54 25.63 16.59 30.36 11.34 13.32 0.65 10.81 9.03
Ada-SKV 24.22 29.92 48.96 40.08 25.55 17.45 31.03 12.12 14.53 0.52 10.57 9.44
Ours. Copy 24.97 30.94 49.45 42.25 26.34 18.54 32.08 12.09 13.88 0.62 10.94 9.38
Ours. Mix 25.59 31.33 50.26 42.66 27.20 19.37 32.74 11.62 15.61 0.50 9.97 9.43
Mistral-7B-Instruct, KV Size=1024
SKV 25.38 30.22 49.29 41.84 26.60 18.08 31.90 11.69 13.89 0.52 10.54 9.16
PyramidKV 24.28 30.05 49.17 40.49 26.43 18.80 31.54 11.77 14.51 0.51 10.19 9.25
Ada-SKV 24.82 31.49 48.80 41.18 27.38 18.22 31.98 11.96 13.82 0.53 9.92 9.06
Ours. Copy 25.87 31.44 49.55 41.95 27.09 19.88 32.63 12.21 14.17 0.50 10.58 9.37
Ours. Mix 25.64 32.54 50.49 41.80 27.88 18.89 32.87 11.94 14.93 0.50 10.49 9.47
Table 4: Details results for different KV cache (64, 128, 256, 512, 1024) on Llama-3-8B-Instruct and Mistral-7B-Instruct.
Appendix B Dataset Details
Table 5 provides details of the datasets used in our experiments. To evaluate the retrieval and contextual reasoning abilities of different KV cache compression methods, we use six question-answering datasets from LongBench (Bai et al., 2024) and four additional question-answering datasets from LooGLE (Li et al., 2024b) as benchmarks. For the LooGLE datasets, we randomly selected 100 examples to form the datasets used in this study.
Source Label Task Task Type Eval metric Avg Len Language Nums
LongBench NrtvQA NarrativeQA Single-Doc. QA F1 18,409 EN 200
LongBench Qasper Qasper Single-Doc. QA F1 3,619 EN 200
LongBench MF-en MultiFieldQA-en Single-Doc. QA F1 4,559 EN 150
LongBench HotpotQA HotpotQA Multi-Doc. QA F1 9,151 EN 200
LongBench 2WikiMQA 2WikiMultihopQA Multi-Doc. QA F1 4,887 EN 200
LongBench Musique Musique Multi-Doc. QA F1 11,214 EN 200
LooGLE Doc.QA Comprehension&reasoning Long Dependency QA F1 15,498 EN 100
LooGLE Info.Retrieval Multiple information retrieval Long Dependency QA F1 14,808 EN 100
LooGLE Timeline Timeline reorder Long Dependency QA F1 15,425 EN 100
LooGLE Computation Computation Long Dependency QA F1 17,001 EN 100
Table 5: Details of Datasets.
Appendix C LongBench Results
We list the evaluation results on the remaining datasets from LongBench (Bai et al., 2024) in Table 6. Our head-level KV cache compression method also outperforms the other baselines, particularly on summarization and Code datasets across Llama-3-8B-Instruct and Mistral-7B-Instruct.
Method Summarization Few-Shot Learning Synthetic Code Avg.
GovReport QMSum MultiNews TREC TriviaQA SAMSum PCount PRe LCC RB-P
Llama-3-8B-Instruct, KV Size=Full
FullKV 28.71 23.26 26.64 73.5 90.48 42.33 4.80 69.25 59.29 54.05 47.23
Llama-3-8B-Instruct, KV Size=128
SnapKV 19.83 21.80 21.41 65.50 89.72 38.71 5.75 69.00 58.74 54.57 44.50
Ada-SnapKV 20.89 22.11 21.68 70.50 90.82 39.20 5.91 69.50 59.75 54.86 45.52
HeadKV-R 21.08 22.35 22.50 71.50 89.45 38.40 5.00 69.50 60.89 59.92 46.06
HeadKV-R2 21.76 22.16 23.94 71.50 90.19 38.88 6.60 69.50 61.08 60.21 46.58
Mistral-7B-Instruct, KV Size=Full
FullKV 32.87 24.24 27.10 71.00 86.23 42.79 2.75 86.98 56.93 54.49 48.54
Mistral-7B-Instruct, KV Size=128
SnapKV 20.76 22.72 21.38 67.00 85.06 40.22 3.51 65.06 52.20 47.01 42.49
Ada-SnapKV 21.13 22.76 22.25 68.50 85.60 41.05 3.33 62.54 52.88 49.25 42.93
HeadKV-R 22.19 22.86 22.57 69.50 85.46 41.16 3.56 74.49 54.60 50.89 44.73
HeadKV-R2 24.30 23.48 24.18 70.50 85.54 40.72 4.83 72.63 55.49 51.39 45.31
Table 6: Results on LongBench (KV cache=128).
Figure 7: Head visualization results on Mistral-7B-Instruct. Our enhanced distribution can also cover the important retrieval heads.
Figure 8: Neele-in-a-Haystack test results on Mistral-7B-Instruct. The results are consist with those on Llama-3-8B-Instruct and our Retrieval-Reasoning Heads distribution (HeadKV-R2 still outperforms other strong baselines.
Appendix D Reasoning-in-a-Haystack details
As shown in Figure 9, we provide examples for each task used in the Reasoning-in-a-Haystack test [^3] https://huggingface.co/datasets/RMT-team/babilong. We use the dataset curated by Kuratov et al. (2024) for the Reason-in-a-Haystack test. Specifically, they use the “Input” listed in Figure 9 as the needle and split it into different sentences using dot. These sentences are then inserted into various positions within the haystack to form the final Reason-in-a-Haystack test examples.
As shown in Figure 10, we provide detailed results across different context lengths and tasks. Our proposed head-level KV cache compression method achieves the best average score among all methods. Notably, on the Llama-3-8B-Instruct model, our method surpasses the Full KV cache on the QA3 task, the most challenging task among the five, as indicated in Kuratov et al. (2024). This suggests that our Retrieval-Reasoning Heads distributions are better at capturing semantically relevant and important content, leading to generating the correct answers.
Figure 9: Examples for each task used in Reasoning-in-a-Haystack test.
Figure 10: Detail results for Reasoning-in-a-Haystack test on five datasets.